Find the transfer function X2/X1 of the given signal flow graph. Use fx to input your answer.Question 8 Find the transfer function X2/X1 of the given signal flow graph. Use fx to input your answer. 40 0 a X10 X2 CD h g
Answer #1

given a d ot g X2 Now Transfer Function = 2 Masons from gain Formula X2 E MK A *\x n=1 Д where K = no. of Forward path Mk ktNow k=2 [m, = acd efh > Forward path gain = are Now Loops L =cb ans t₂ = fg A = 1- [cb+fg] A = 1-2b-f9 A 1 of- fg - A,= =/fgX₂ T = Х acd[1-49]+ eth[i-ch] 1-cb-fg acdf2= +9]+eth [l-c6] I-cb-fg X₂ = Х
Answer #2
5. Find out transfer function of control system by using Signal flow Graph method, given below in figure 2
source: H
answered by: Viral